carmelle:

What are the values a, b, and c in the following quadratic equation? 18 = −9x + 7x^2

jhonyy9:

make this form ax^2 +bx +c = 0

carmelle:

ohh 7x^2 -9x + 18 = 0

jhonyy9:

not plus 18 attention !!!

carmelle:

oh

carmelle:

7x^2 -9x - 18 = 0

jhonyy9:

exactly

carmelle:

Thank you!!

jhonyy9:

np yw